The United States government claims DUI arrests average 22,096 per state per year. A sample of size n=10 states finds the mean to be 44,002. If DUI arrests are normally distributed with a standard deviation of 28,584. Develop null and alternative hypothesis that will help deciding whether DUI arrests are higher than government claims at the 1% significance level and calculate the p value
Answer #1

Calculated t value = xbar-mu / SD/(n)1/2

= 44002-22096/28584/(10)1/2

= 2.423

t value as calculated from the table for degree of freedom (n-1=9) with 0.01 level of significance.

t=2.821

Since the computed value is less than the standard value, the null hypothesis is accepted.

p value is 0.0192